Set up an equation and solve: the area of a rectangle is 40 square feet. the length is 2 feet less than twice the width. find the dimensions.
Answer by TimothyLamb(4379)

Lw = 40
L = 2w - 2
---
Lw = 40
(2w - 2)w = 40
2w^2 - 2w - 40 = 0
---
the above quadratic equation is in standard form, with a=2, b=-2, and c=-40
---
to solve the quadratic equation, plug this:
2 -2 -40
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
w = 5
w = -4
---
negative width doesn't make sense for this problem, so use the positive root:
w = 5
L = 8
---
Answer:
the speed of a train: 72 km / 1.5 hours = 48 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php